Suppose equation of motion for one dimensional oscillator is given by: ?̈ + ??̇ + 9?...

  1. Suppose equation of motion for one dimensional oscillator is given by:

    ?̈ + ??̇ + 9? = 0

    1. For α values of 3, 6, and 9 indicate what kind of oscillatory system it would be.

    2. Find expression for x(t) for each value with the initial conditions, x0 = 0 and vo = 5 m/s.

      Use proper ansatz to start from scratch (Check whether these initial conditions might

      be non-sense. Choose convenient initial conditions whenever necessary).

    3. Solve the equation with α = 6 and with a drive force of f(t) = 2 cos(3t)
Answer #1

d 2x The tada JE +92 = 0 2C 10) = 0 wcom - xro) = 5 mis Taking a=3 022 + 3d2c +9X1V=O dt2 dt char-eqa m²+3m+g (m+ 2)² + 27 mXlt = (6+czt) et 10) = 0 using Scio) - Tazo system is Non- oscillatory. xilo) = 5 * C2+ -3C = 5 J C2 = 5 x 9 = 5+ E3F For od==(atrat) et For x=6 f=26053 XIH = act Xp xp = 2 COS3 2 Cost (M43)2 missmislu ap - m (2 Cos31 1 m²-9 6 m2 Scp 1m²-g 2 Cost 6m

Hope it helps you out.

In case any doubt please feel free to ask.

I will be happy to help you.

Thanks!!